If TC= 3+Q^2 and demand Is P=3-1÷2Q. Calculate total profit
1
Expert's answer
2021-07-26T11:00:01-0400

Using the cost function, marginal cost would be:

"\frac {dTC} {dQ} =MC=2Q"

Using the value of MC, market quantity will be the one at which MC equates to price such that:

"MC=P"

"2Q=3-(\frac {1} {2} )Q"

"2Q+(\frac {1} {2} )Q=3"

"(\frac {5} {2} )Q=3"

"Q=\frac {6} {5}"

At the quantity of 6/5 units, market price would be:

"P=3-(\frac {1} {2} )(\frac {6} {5} )"

"P=3-(\frac {3} {5} )"

"P=\frac {12} {5}"

Therefore, at the price of "\frac {12} {5}" and quantity of "\frac {6} {5}" firms profit would be=

"=TR-TC"

"=Q[3-(\frac {1} {2} )Q]-(3+Q^2)"

"=(\frac {6} {5} )[3-(\frac {1} {2} )2]-[3+(6)5)^2)]"

"=[(\frac {6} {5} )×2]-[3+1.44]"

"=-2"

Therefore, firm will incur loss.